# Neuromorphic Overparameterisation: Generalisation and Few-Shot Learning in Multilayer Physical Neural Networks

### v4

@@ Acknowledgements
Authors patent applicant. Inventors (in no specific order): Kilian D. Stenning, Jack C. Gartside, Alex Vanstone, Holly H. Holder, Will R. Branford. Application number: PCT/GB2022/052501. Application filed. Patent covers the method of programming deep networks.
Methods
Some description of our methodologies are reproduced from earlier work of several of the authors[61, 5], as similar methods are employed here. The methods section is organised as follows: Experimental Methods includes the fabrication of samples, measurement of FMR response and the details of implementing reservoir computing and interconnecting arrays. Following this we discuss the tasks chosen and how they are evaluated in Task selection. Learning algorithms then provides a detailed description of the learning algorithms used in this work.
Experimental methods
Nanofabrication
Artificial spin reservoirs were fabricated via electron-beam lithography liftoff method on a Raith eLine system with PMMA resist. 25 nm Ni81Fe19 (permalloy) was thermally evaporated and capped with 5 nm Al2O3. For WM, a ‘staircase’ subset of bars was increased in width to reduce its coercive field relative to the thin subset, allowing independent subset reversal via global field. For PW, a variation in widths were fabricated across the sample by varying the electron beam lithography dose. Within a 100 m 100 m write-field, the bar dimensions remain constant. The flip-chip FMR measurements require mm-scale nanostructure arrays. Each sample has dimensions of roughly 3x2 mm. As such, the distribution of nanofabrication imperfections termed ‘quenched disorder’ is of greater magnitude here than typically observed in studies on smaller artificial spin systems, typically employing 10-100 micron-scale arrays. The chief consequence of this is that the Gaussian spread of coercive fields is over a few mT for each bar subset. Smaller artificial spin reservoir arrays have narrower coercive field distributions, with the only consequence being that optimal applied field ranges for reservoir computation input will be scaled across a corresponding narrower field range, not an issue for typical 0.1 mT or better field resolution of modern magnet systems.
MFM measurement
Magnetic force micrographs were produced on a Dimension 3100 using commercially available normal-moment MFM tips.
FMR measurement
Ferromagnetic resonance spectra were measured using a NanOsc Instruments cryoFMR in a Quantum Design Physical Properties Measurement System. Broadband FMR measurements were carried out on large area samples mounted flip-chip style on a coplanar waveguide. The waveguide was connected to a microwave generator, coupling RF magnetic fields to the sample. The output from waveguide was rectified using an RF-diode detector. Measurements were done in fixed in-plane field while the RF frequency was swept in 20 MHz steps. The DC field was then modulated at 490 Hz with a 0.48 mT RMS field and the diode voltage response measured via lock-in. The experimental spectra show the derivative output of the microwave signal as a function of field and frequency. The normalised differential spectra are displayed as false-colour images with symmetric log colour scale[61].
Data input and readout
Reservoir computing schemes consist of three layers: an input layer, a ‘hidden’ reservoir layer, and an output layer corresponding to globally-applied fields, the nanomagnetic reservoir and the FMR response respectively. For all tasks, the inputs were linearly mapped to a field range spanning 35 - 42 mT for MS, 18 - 23.5 mT for WM and 30-50 mT for PW, with the mapped field value corresponding to the maximum field of a minor loop applied to the system. In other words, for a single data point, we apply a field at + then -. After each minor loop, the FMR response is measured at the applied field - between 8 - 12.5 GHz, 5 - 10.5 GHz and 5 - 10.5 GHz in 20 MHz steps for MS, WM and PW respectively. The FMR output is smoothed in frequency by applying a low-pass filter to reduce noise. Eliminating noise improves computational performance[5]. For each input data-point of the external signal , we measure distinct frequency channels and take each channel as an output. This process is repeated for the entire dataset with training and prediction performed offline.
Interconnecting arrays
When interconnecting arrays, we first input the original Mackey-Glass or sinusoidal input into the first array via the input and readout method previously described. We then analyse the memory and non-linearity of each individual frequency output channel (described later). A particular frequency channel of interest is converted to an appropriate field range. The resulting field sequence then applied to the next array via the computing scheme previously described. This process is then repeated for the next array in the network. The outputs from every network layer are concatenated for learning.
Task selection
Throughout this manuscript, we focused on temporally-driven regression tasks that require memory and non-linearity. Considering a sequence of T inputs , the physical system response is a series of observations across time. These observations can be gathered from a single reservoir configuration as in Figure 1, or can be a collection of activities from multiple reservoirs, in parallel or inter-connected as in Figure 2. In other words, the response of the system at time t is the concatenation of the outputs of the different reservoirs used in the architecture considered. The tasks faced can be divided into five categories:
Sine Transformation tasks. The system is driven by a sinusoidal periodic input and asked to predict different transformations, such as . The inputs are chosen to have 30 data points per period of the sine, thus with . The total dataset size is 250 data points. If the target is symmetric with respect to the input, the task only requires nonlinearity. If the target is asymmetric, then both nonlinearity and memory are required.
Mackey-Glass forecasting. The Mackey-Glass time-delay differential equation takes the form and is evaluated numerically with = 0.2, n = 10 and = 17. Given as external varying input, the desired outputs are , corresponding to the future of the driving signal at different times. We used 22 data points per period of the external signal for a total of 250 data points. This task predominantly requires memory as constructing future steps requires knowledge of the previous behaviour of the input signal.
NARMA tasks. Non-linear auto-regressive moving average (NARMA) is a typical benchmark used by the reservoir computing community. The definition of the x-th desired output is , where the constants were set to A = 0.3, B = 0.01, C = 2, D = 0.1. The input signal used is the Mackey-Glass signal, where the variable is introduced to account for a possible temporal shift of the input. For , is the application of NARMA on the Mackey-Glass signal at the current time t, while for , is the result of NARMA on the input signal delayed by ten time steps in the future. The index can instead vary between one, defining a task with a single temporal dependency, and fifteen, for a problem that requires memory of fifteen inputs. Varying and , we can define a rich variety of tasks with different computational complexity.
Evaluations of memory-capacity (MC) and non-linearity (NL). Memory capacity and non-linearity are metrics frequently used for the characterization of the properties of a physical device. While these metrics do not constitute tasks in the common terminology, we include them in this section for simplicity of explanation. Indeed, we used the same training methodology to measure MC and NL as in the other tasks faced. We evaluate these metrics with the Mackey-Glass time-delay differential equation as input. This gives results that are correlated to conventional memory-capacity and nonlinearity scores, with some small convolution of the input signal - negligible for our purposes of relatively assessing artificial spin reservoirs and designing network interconnections.
For memory-capacity the desired outputs are , corresponding to the previous inputs of the driving signal at different times. To avoid effects from the periodicity of the input signal, we set k=8. The R2 value of the predicted and target values is evaluated for each value of where a high R2 value means a good linear fit and high memory and a low R2 value means a poor fit a low memory. The final memory-capacity value is the sum of R2 from = 0 to = 8.
Memory capacity and non-linearity can be calculated using a single multiple frequency channels. For the single channel analysis, we perform the same calculations, but using just a single FMR channel. -
Frequency decomposition, a few-shot learning task. The network is driven by a sinusoidal input and needs to reconstruct a decomposition of a temporal varying signal in the form of , where . The values of and are randomly sampled at the beginning of each task from uniform distributions. In particular, and . The output layer is composed of nodes, and the system is asked to predict a target after observing the values of over K data points i.e. time steps. The value of K adopted for the examples of Figure 4 b) and c) is fifteen, but the network reports good performance even with (Figure 4 d) and e). To face this challenging task, we used the PNN in the overparameterised regime and a meta-learning algorithm to quickly adapt the read-out connectivity. Details of the meta-learning approach are given below in the Meta-learning section of the Methods.
Learning algorithms
The type of learning algorithm used to select features and train the networks varies throughout the manuscript.
For comparisons between different systems (Figures 1 2), training is accomplished through a features selection algorithm (discussed below) and optimisation of the read-out weights . When exploring the effects of overparameterisation, we randomly select features and then optimise for those features.
We call the representation at which the read-out weights operate, and we define the output of the system as . The vector simply contains a subset of features of defined via the feature selection algorithm or randomly picked as in the dimensional study on overparameterisation. In this setting, optimisation of is achieved with linear or ridge regression, which minimises the error function .
Feature selection
[…]

### v5

@@ Acknowledgements
Authors patent applicant. Inventors (in no specific order): Kilian D. Stenning, Jack C. Gartside, Alex Vanstone, Holly H. Holder, Will R. Branford. Application number: PCT/GB2022/052501. Application filed. Patent filed in UK through Imperial College London. Patent covers the method of programming deep networks. The remaining authors declare no competing interests.
Supplementary Information
Supplementary note 1 - Ferromagnetic resonance spectroscopy and magnetic force microscopy images of the artificial spin systems
The nanomagnetic arrays in this work are based on square and pinwheel artificial spin ice[66]. Supplementary Figure 1 shows scanning electron micrographs (SEM) (a,c,e) and ferromagnetic resonance (FMR) spectra after AC-demagnetisation (b,d,f). Three arrays were fabricated:
MS is a square artificial spin ice (Supplementary Figure 1 a,b). Bars are high aspect-ratio (530 nm 120 nm) and only support macrospin states[5]. WM is a width-modified artificial spin-vortex ice with a subset of wider, lower-coercivity bars (Supplementary Figure 1 c,d). Bars are 600 nm 200 nm (wide-bar)/125 nm (thin-bar). Wide bars host both macrospin and vortex states[5] whereas thin bars host just macrospins. PW is a pinwheel-lattice artificial spin-vortex ice (Supplementary Figure 1 e,f)[67, 68] with higher density and inter-island coupling. A gradient of bar dimensions are patterned across the sample, ranging from fully-disconnected (e, top) to partially-connected islands (e, bottom) giving a complex range of spectral-responses (Supplementary Figure 1 f). Bar dimensions are constant across 100 2 (length 450 nm, width 240 nm / 265 nm (lower / upper panel) in Supplementary Figure 1 e). Islands support macrospins and vortices.
MS FMR spectra comprise two dominant modes at 9.5 and 12 GHz with opposite linear gradients corresponding to macrospins aligned parallel (positive gradient) and anti-parallel (negative) to Happ (Supplementary Figure 1 b). In both the field-saturated (Supplementary Figure 1 c) and AC-demganetised state (Supplementary Figure 1 d), only macrospin states are observed due to the high aspect-ratio of these bars.
The WM FMR spectra comprises four dominant modes with rich responses: wide and thin bar linear macrospin modes (7 GHz and 9 GHz respectively), -shaped vortex mode (2-6 GHz) and ‘whispering-gallery’-like high-frequency vortex mode[69] (10 GHz) (Supplementary Figure 1 f). Vortices are observed in the wide-bars when the sample is in an AC-demagnetised state.
PW displays a highly complex and non-linear FMR spectra (Supplementary Figure 1 j) due to the broad range of bar dimensions and state variations in this sample. Bars host both macrospins and vortices (Supplementary Figure 1 k,l)
Supplementary Figure 2 shows scanning electron micrographs of variation of dimensions across the MS (a,b) and PW (c-f) samples. In MS, two separate regions with widths of 124 nm and 137 nm are present. This manifests as slightly different coercive fields and resonant frequencies. In PW, a gradient of sample dimensions are fabricated across the surface of the chip. Bar subsets have slightly different widths to further enhance sample complexity. Widths range from 212 - 226 nm (thin) and 243 - 267 nm (wide) giving a broad distribution of resonant frequencies and coercive fields, enhancing output nonlinearity. Furthermore, some bars are connected (panel f) giving rise to complex magnetisation profiles and FMR spectral evolution.
Supplementary note 2 - kmax selection for metric calculation.
Memory-capacity and nonlinearity are typically obtained by inputting a random input signal, with no correlation between consecutive inputs, to the reservoir. When calculating memory-capacity, this ensures that any observed memory effects arise from reservoir states alone. However, our measurement scheme precludes this method. Data is encoded in magnetic field amplitude and readout in-field resulting in field-dependent frequency shifts in the FMR response. For random inputs, the sharp jumps between consecutive inputs causes information to shift between outputs (e.g. shifts of 0.5 GHz spanning 25 output frequencies are observed over a 5 mT input range). Linear regression is unable to process this type of shifting leading to misleadingly low calculated memory-capacity.
As such, we use the smoothly varying Mackey-Glass equation as an input signal to calculate the metrics. This ensures that any field shifts between consecutive input are minimised. Our FMR peaks are broad with microstate information held across multiple neighbouring outputs (e.g. 7 GHz and 7.02 GHz will be collinear). A disadvantage of this approach is that the Mackey-Glass equation is quasi-periodic. In addition to memory from the artificial spin reservoir states, this can lead to a memory arising from the similarity between current and previous inputs. As such, the value of kmax with which the metrics are calculated must be chosen carefully to minimise the effects of the periodic input.
Supplementary Figure 3 shows the memory-capacity and nonlinearity of the three artificial spin reservoir samples. Memory-capacity of the input signal with itself is also shown. All curves show characteristic humps and continuous rising in the memory-capacity profile due to the periodic nature of the MG input. Nonlinearity shows no variation with kmax. We chose kmax = 8 as this is at the end of the flat region on the input memory-capacity curve i.e. this is the maximum value before self-correlation from the next period starts.
Supplementary note 3 - Sine and NARMA transforms for each artificial spin reservoir
Supplementary Figure 4 a) shows the MSE for each reservoir when transforming a sinusoidal input to a variety of targets. Signal-transformations require varying memory-capacity and nonlinearity depending on the target waveform. Targets which are symmetric w.r.t. the input (eg. sin2(x), sin(3x)sin(x)) require nonlinearity only, asymmetric waveforms w.r.t the input (e.g. saw(x), sin(x/2)) require both nonlinearity and memory-capacity as equivalent input values must be transformed to different output values across the wave-cycle. Highest nonlinearity PW dominates for symmetric transforms. Highest memory-capacity WM dominates for asymmetric transforms. Example plots for asymmetric |sin(x/2)| transform are shown in Supplementary Figure 4 c).
Supplementary Figure 4b) shows performance when performing a NARMA-transform[41] on the Mackey-Glass input, evaluated as with A = 0.3, B = 0.01, C = 2, D = 0.1. X varies between 1-15, defined as ‘NARMAX’. This task adds additional nonlinear complexity vs. normal Mackey-Glass prediction, favouring both memory-capacity and nonlinearity with WM and PW performing similarly across all tasks. Example plots for the most-challenging NARMA7 are shown in Supplementary Figure 4 d).
Supplementary note 4 - Per-channel metric analysis for series connections
We assess metrics on a per-output channel basis, highlighting that memory and nonlinearity are provided by distinct spectral channels. Figures 5 a-c) (top) show FMR spectra at max (dark blue) and min (light blue) input field-amplitude, correlation to previous time steps for 0-7 previous time-steps (n) where Corrn is the correlation to the nth previous input, total correlation and nonlinearity of each output frequency-channel (bottom). Here, the correlation calculation is identical to the memory-capacity calculation, except that only one channel is used rather than multiple channels. As such, we call it correlation to avoid confusion with other discussions of memory capacity. For MS, high Corrn is limited to n 3. In contrast, WM and PW have some outputs which are correlated to 0-3 prior time-steps and others correlated to 4-7 steps (e.g. WM, 7.2 - 7.5 GHz). The presence of multiple correlation timescales is provided by vortex dynamics and is key to the strong prediction performance observed later. In PW, the main FMR mode has high nonlinearity due to complex disordered microstate dynamics. The gradient of physical structures in the array provide more non-degenerate nonlinear responses and hence the highest nonlinearity score.
Supplementary note 5 - Vortex induced memory amplification.
Supplementary Figures 6 a-c) show the memory-capacity of reservoir 1 (MCin) vs memory-capacity of reservoir 2 (MCout)=when R2 is a) MS, b) WM and c) PW. Supplementary Figures 6 d-f) shows the relationship for MCin vs MCout when calculating memory-capacity for specific previous inputs. In all cases, a linear trend is observed. For MS, the gradient of MCin vs MCout stays approximately the same as memory-capacity is evaluated to further previous inputs. For WM and PW, strong MCout is observed throughout. The gradient between input and output memory-capacity increases for further previous inputs. As such, the memory of the input signal is effectively amplified. The effect is strongest for WM. Small hints of long-term memory in the input signal translate to large contributions in the second reservoir. This provides additional information into why the ordering of reservoirs is crucial, as memory amplification at the end of the network is key for improved performance.
Supplementary note 6 - Interconnecting high-dimensional physical systems
No individual physical system excels across all tasks, and any single system tends to underperform for harder tasks. This is a well-known symptom of single reservoir systems in software and hardware[70, 71, 26, 38, 20]. In software, multiple reservoirs with distinct responses have been combined into networks to harness the benefits of different dynamical behaviours[70, 71, 26, 38, 20]. In such networks, each reservoir can be viewed as a complex node[72] with high-dimensionality and distinct memory and nonlinearity scores.
Here, we construct parallel and series (often termed hierarchical or deep) networks of physical reservoirs, combining a synergistic suite of distinct nanomagnetic arrays with substantially enhanced performance. The increased output dimensionality gained from multiple physical arrays is beneficial for computation, but increases the likelihood of overfitting - a common challenge in machine-learning. One way to avoid this is to increase the size of the training dataset, however, this luxury is often unavailable in real-world applications, especially in remote physical use-cases which must expensively acquire their own data using sensors. In response, we implement a feature selection scheme [20, 42] to reduce the number of network outputs by discarding less useful channels, avoiding overfitting and providing robust, accurate performance (see Methods).
Schematics of the parallel and series networks are shown in Supplementary Figure 7 a) and b) respectively. In parallel networks, data is input into multiple arrays and the FMR response of each array measured. During offline training, the response of different arrays is concatenated to give the network output (i.e. two arrays with 300 frequency output channels produce a network with 600 outputs/parameters).
In series networks, data is input to the first array/reservoir (R1) and FMR response is measured. The 300-dimensional R1 FMR response now must be converted to a 1D field input for the second reservoir array R2. To accomplish this, we take the amplitude of a specific FMR frequency-channel and map it to an input field sequence for R2. The R1 output frequency-channel is selected via per-channel memory-capacity and nonlinearity evaluation. This analysis and scaling step is performed offline, with future prospects for on-chip implementations. The FMR spectra of R1 and R2 are then combined for offline training and prediction[20].
Parallel networks are capable of enhancing performance at tasks demanding high nonlinearity and low memory. Example transformations are shown in Supplementary Figure 7 e) for sin(3x) + cos(x) and |sin(x/2)| where improvements up to 4.4 vs. the best single array observed when combining all three arrays in parallel (further transforms are discussed later).
Supplementary Figure 7 c,d) shows series network MSE vs for future prediction (c) and NARMA-transformation (d) of Mackey-Glass time-series. The best single reservoir (WM, black) parallel network (PW + WM + MS, brown) are also shown. Parallel networks do not show significant enhancement due to the lack of memory-transfer in this architecture. Low memory-capacity (R1) to high memory-capacity (R2/R3) series network architectures significantly improve performance, up to 2.7 and 4 in three-layer deep networks for challenging t+7 and NARMA7 tasks respectively with predictions shown in Supplementary Figure 7 i). As in the brain, reservoir ordering is critical[20], with high memory-capacity to low memory-capacity network architectures showing weak improvements of 1.5 for all prediction tasks (supplementary note 7 discusses deep architecture ordering). Three-layer deep networks substantially outperform two layers, highlighting the benefits of more complex networks.
The best prediction is found for MS -> PW -> WM (3 layers, memory-capacity = 6.7, nonlinearity = 0.75). While the 3-layer deep network has a memory-capacity improvement of 1.31 vs. the best single reservoir, MSE values are up to 3.7 lower than the single WM sample for future prediction tasks due to the expanded range of temporal responses/timescales in the deep network. This expanded range of temporal responses results in strong flattening of MSE periodicity vs. single and parallel reservoirs, especially evident in the NARMA-transform task (f). This enhanced temporal richness is not accurately reflected in a single memory-capacity/memory-capacity score. Memory-capacity and nonlinearity metrics are valuable guides for basic reservoir evaluation, but lack the granularity for accurately predicting performance.
Supplementary Figure 8 a) shows MSEs of the best single and parallel networks for each signal-transformation task. Only the best 2- and 3-layer parallel architectures are shown. Lower MSE is observed for parallel architectures for asymmetric tasks requiring memory-capacity and nonlinearity. Parallel networks are able to harness memory-capacity and nonlinearity characteristics from different artificial spin reservoirs to reduce MSE for asymmetric transforms. Lowest MSEs are observed when combining WM + PW, or all three artificial spin reservoirs with performance gains up to 4.4 for more complex asymmetric tasks. For some tasks, adding MS to create a 3 layered parallel architecture worsens performance due to the limited transformation capabilities of that reservoir replacing more useful outputs from the other artificial spin reservoirs.
Interestingly, the best parallel architectures have lower nonlinearity than the single PW (0.65 vs. 0.75). The feature selection process optimises outputs across all tasks. Introducing high memory-capacity outputs from WM to improve asymmetric tasks comes at the cost of sacrificing high nonlinearity outputs from PW, reducing nonlinearity-only task performance.
Supplementary Figure 8 b) shows MSE values for single, parallel and series networks for signal-transformation. Series networks show enhanced performance for complex asymmetric tasks requiring both MC and NL (e.g. sin(2x) + cos(3x)). For simpler symmetric and asymmetric tasks, single and parallel networks dominate.
Supplementary Figure 9 shows the future prediction and NARMA transformation performance in a deep network when going from a reservoir with high memory-capacity to a reservoir to low memory-capacity. Low improvements are observed throughout. The initial reservoir response obscures short term information. The second reservoir simply mimics this response. As such, short term information is lost.
The ordering of artificial spin reservoir’s plays an important role: if R1 has low memory-capacity, its response is captures short term behaviour. When this information is transferred, a high memory-capacity R2 receives information about short-term behaviour which it can retain for longer. As such, short-term memory is retained in R1 and long-term memory is retained in R2. For high-to-low memory-capacity, R1’s output will be obscured by history-dependence, limiting the amount of information retained about short term behaviour. When this information is transferred, R2 simply mimics this information. As such, neither R1 or R2 retain a reasonable short-term memory, only long-term correlations are present, reducing the overall memory-capacity.
Complex memory-capacity + nonlinearity signal-transformations require higher harmonic generation and responses shifted in phase w.r.t the input. By going from a high memory-capacity reservoir (WM) to a high nonlinearity reservoir (nonlinearity), both a linear and nonlinear representation of the input signal is present in the overall network (linear from high memory-capacity, nonlinear from high nonlinearity) producing a diverse set of lagged responses and higher harmonics.
Supplementary note 7 - Mackey-Glass attractor reconstruction
We use another methodology to understand the efficacy of our method in reconstructing the original time-series, even in the case of the prediction task. In Supplementary Figure 10 we plot show two prediction tasks (t+1 and t+7) for the Mackey-Glass time series (top curves). We see that at low prediction windows, both a single reservoir and a network of reservoirs reconstruct the time series. We can analyse the efficacy of the task by looking at the reconstructed attractors using the State Space Reconstructor method [73, 74, 75], embedding it in 3-dimensions with a delay of . We compare the original reconstructed attractor (black curves) versus the single (green curves) and PNN models (red curves). These plots contain much more information than MSE (center middle plots). As we can see, at low prediction windows (t+1) the reconstructed attractor is well approximated by both the RC models. However, at longer prediction windows (t+7) the PNN follows more closely the attractor. This can be quantified by looking at the multi-dimensional variance of the 3-dimensional time series obtained by subtracting the trajectory of the attractor from the original time-series versus the two reconstructed ones (bottom plots). We see error is much more spread at longer prediction windows for the single RC.
Supplementary note 8 - PNN in the underparameterised regime
In Figure 2 of the main text, the featuer selection algorithm is allowed to freely chose between all outputs. The total number of outputs used for each architecture is: WM - 34, MS+PW+WM - 44, MSWM - 42, MSWMPW - 93 and PNN - 13567. The enriched readout dimensionality is a distinct advantage of the PNN in terms of computational performance.
Supplementary Figure 11 shows the performance when the PNN is constrained to have a similar number of outputs the other networks. Here, 97 PNN are used during training. We find that performance is at a similar level to the three-series network. Performance is worse for some tasks. This is because the feature selection algorithm is stochastic and requires more iterations to explore the entire readout space. Additionally, the algorithm finds a set of outputs that performs well at all tasks. Given enough iterations, the performance of the PNN will at least match other architectures, as the PNN contains outputs from all other network architectures.
Supplementary note 9 - Software comparison
We now compare the performance of the various network architectures to software models, specifically echo state networks (ESN) and multilayer perceptrons (MLP). For ESN, we vary the number of internal nodes (Nnodes) and for each node we randomly initialise 50 networks as described in the Methods section ESN Comparison. For MLPs, we vary the size of the hidden layers (Nhidden) from 1-500 and number of previous inputs the MLP receives (Tseq) from 1-10 (see Methods for MLP details).
Supplementary Figure 12 shows the performance of ESNs with varying number of nodes (Nnodes) for a) Mackey-Glass future prediction, b) NARMA transform, c) NARMA transform plus future prediction as presented in the main text. Panels d-f) show the ESN performance when varying Ntrain and Nnodes for each task. In a-c), the blue line represents the average error of 50 randomly initialised ESNs for a given number of internal nodes with shaded area representing . Coloured lines show the performance of the various networks explored in the main text. In a-c), Ntrain = 200. MSE represents the average performance over multiple tasks: t+0 to t+12 for future prediction (a), NARMA 0 - 12 for NARMA transforms (b) and t+0 to t+20 for NARMA plus future prediction (c). For ESNs, as Nnodes and Ntrain increases, MSE reduces as expected. For each task, we find that single arrays are matched by ESNs with 20 - 40 nodes. As the physical network complexity increases, the corresponding ESN size to match performance also increases. The ESN size required to match the PNN performance is 100 for the NARMA task and 500 for prediction tasks.
We now compare the physical network performance to MLPs. MLPs are static, i.e. they hold no information about previous inputs and will fail at predictive tasks. As such, we vary the number of previous inputs given to the MLP during each time step. Supplementary Figure 13 shows the performance of MLPs when varying the size of the hidden layers (Nhidden) and the number of previous inputs provided to the MLP (Tseg) for a) Mackey-Glass future prediction, b) NARMA transform, c) NARMA transform plus future prediction as presented in the main text. As with the ESN, MSE represents the average performance over multiple tasks: t+0 to t+12 for future prediction (a), NARMA 0 - 12 for NARMA transforms (b) and t+0 to t+20 for NARMA plus future prediction (c). The MLP MSE reduces as Nhidden and Tseg increases as expected. For Tseg = 1, the physical networks outperform all trialled MLPs due to the lack of past information in the MLP. We find that series networks are well matched to MLPs with Tseg = 2 and Nhidden 10 and the PNN is matched to an MLP with Tseg = 3 and Nhidden 10-50 for all tasks. Whilst the resulting MLP is simple, the network interconnections must be trained using gradient descent. This is fairly inexpensive for small networks, but as network and task complexity increases, so does the expense of training MLP weights.
We now implement the same networking methodology described in the main text with software echo-state networks.
We begin by initializing ESNs with 200 nodes (to ensure a similar number of output channels in comparison to nanomagnetic arrays) and select three which display similar memory-capacity and non-linearity to the three nanomagnetic arrays. Supplementary Figure 14 shows the memory-capacity and non-linearity of the three nanomagnetic arrays alongside three ESNs with similar properties (red circles). The ESNs were found via randomly varying the hyperparameters of the ESN and selecting ESNs which had the closest characteristics.
We find that, despite the ESNs and nanomagnetic arrays displaying similar metrics, the performance shows different characteristics. For the two high memory-capacity ESNs, the MSE rises slightly from future step 1 – 3, and then becomes flat, whereas the hardware reservoirs display a periodic profile in the MSE. We find that it is challenging to find an ESN that displays the same characteristics as the hardware reservoirs. This is not surprising, as nanomagnetic reservoirs and software ESNs are governed by different underlying dynamics. In nanomagnetic arrays, coupling arises from the dipolar field which decays spatially and hence coupling only occurs locally, whereas in ESNs, coupling is randomly assigned across the network. In ESNs, the coupling strengths between nodes are initialised and remain fixed, whereas in nanomagnetic arrays, the coupling between neighbouring elements depends on the state of those elements (e.g. uniform magnetisation or vortex). Furthermore, in ESNs, only certain nodes are coupled to the input, which enhances memory-capacity as certain nodes require a number of network updates before receiving information about a particular input. On the other hand, in nanomagnetic arrays, each node is subject to the input data when it is first applied. Finally, experimental nanomagnetic systems are subject to noise, whereas ESNs are noise free. It is possible that, with extensive optimisation, an ESN could be initialised that has more similar characteristics to nanomagnetic arrays, however, achieving this is beyond the scope of this report.
We note that when initialising ESNs, the majority of them had memory-capacity far beyond the nanomagnetic arrays (close to 8).
We now interconnect these arrays following the same methodology used to connect different nanomagnetic arrays. We begin by analysing the channel specific time correlation and non-linearity of the three ESNs displayed in Supplementary Figure 15.
When comparing these results to the nanomagnetic reservoirs (SI Supplementary Figure 5) we again see fundamental differences in the dynamics of ESNs vs nanomagnetic hardware. For the low memory-capacity ESN (ESN 1), we find that all channels are only correlated with short-term previous inputs. Conversely, for the higher memory-capacity ESNs (ESN 23), there are a broad range of correlations. Some channels are correlated with short term previous inputs, whereas some are correlated with previous inputs many time-steps ago. In contrast, in nanomagnetic arrays, all of channels are most strongly correlated with short-term inputs and only a handful are correlated with long-term previous inputs. This is a symptom of the differences between internode couplings in ESN when compared to nanomagnetic arrays. The random, sparse connections in ESN allow many different dynamic timescales whereas the spatially constricted, local couplings in nanomagnetic systems prevent this behaviour.
For each ESN, we now select six channels: three with the highest memory capacity and three with the highest non-linearity, giving a total of 18 input time-series to feed into the next layer. Each ESN is the subject to all input sequences to produce 54 2-series networks. We then create 3-series networks by selecting the 2-series architectures which: 1) begin with the low memory-capacity reservoir (ESN 1) and 2) which display the lowest MSE. We then perform the same per-channel metric analysis, generating 9 input sequences which are passed to the final ESN to form a 3-series network. Finally, we combine the responses from all single, 2-series and 3-series networks using the same methodology as the PNN. We evaluate the performance of the ESN networks using the same feature selection methodology described in the Methods section of the manuscript.
Supplementary Figure 16 compares the performance of single, 2-series, 3-series and PNN architectures for the three prediction tasks presented in the main text. The top row shows the nanomagnet response previously included in the manuscript and the bottom row shows the ESN network results. The ESN results displayed are taken from the architecture with the lowest MSE over all future / NARMA steps for a given task. When interconnecting ESNs, we find that whilst improvements are observed for some tasks, they are less pronounced than for the physical system. Series networks only show improvements for future prediction t+1 and t+2, NARMA 10 and above, and NARMA + Future prediction t+4 to t+13, whereas series networks improve all predictions for the nanomagnetic arrays. Software analogues of the PNN architecture show improvements for more tasks (notable future prediction t+7 – t+12), but also have worse performance for other tasks. We believe this is because single ESNs already display a rich set of memory time-scales, and hence interconnecting arrays does not enrich the overall network to the same extent. These results give us an indication that interconnecting physical systems can give MSE improvements which produce a similar level of performance to software counterparts. However, we stress here that the initial ESNs selected are not optimised for performance. If we were to optimise ESNs for performance, then it is likely that the software network would outperform nanomagnetic hardware.
In addition, we test deep echo state networks, where multiple interconnections between different ESN layers are made (i.e. multiple nodes from one ESN connect to multiple nodes in another) in a similar manner to previous work of some of our authors[20]. We connect the three ESNs with similar characteristics to the physical nanomagnetic reservoirs into a 3 x 3 network, whereby each ESN is present in each layer. Layer interconnections are initialised randomly and we average over 10 random trials. The prediction results for the three tasks are shown in Supplementary Figure 16 (red curve, ‘Deep ESN’). As expected, the performance dramatically improves compared to our proposed method as there is an increase in the amount of information being transferred between different layers. Currently, such interconnections are not possible with the hardware approach as each input is only able to accept one input per time step. If a physical system is used which can take multiple inputs at each time step (e.g. memristors), we can connect it in this way and achieve a computationally powerful network with just a handful of nodes.
Supplementary note 10 - Further overparameterisation details
We now provide further details regarding the overparameterisation regime, specifically, how train length, network sub components and task affect the ability to reach a beneficial overparamterised regime.
We begin with a discussion of how the network subcomponents and architecture affect the effective dimensionality of the output and therefore ability to reach a beneficial overparameterised regime. Supplementary Figure 17 shows the dependence of the number of trainable parameters (i.e. number of output channels used during regression) on the MSE for a) single, b) parallel and c) 2-series networks. Here, MSE is an average over 6 prediction tasks (see methods for details) encompassing tasks that all networks perform well and those which simpler networks struggle. Beginning with the single arrays, we see that no single array benefits from being in an overparameterised regime as the minimum MSE in this region is higher than the minimum MSE in the underparameterised regime. However, we see that the underlying array affects the ratio of MSEOP to MSEUP. PW shows a far greater reduction in the overparameterised regime compared to WM and MS. Combining these results with the memory capacity and non-linearity calculations in and the spectral evolution during computation in Figure 1 of the main text, we see that having a diverse spectral output is key to achieving a beneficial overparameterised regime. PW has structural diversity, producing a large variety of magnetic states which are detected at different resonant frequencies. MS is a highly linear reservoir, with relatively simple dynamics and magnetic elements which all occupy the same frequency space. WM is a linear sample with strong memory capacity, with magnetic elements largely occupying the same frequency space, but behaving in a complex manner. For an equivalent frequency range and resolution, PW fills out the frequency space, producing a diverse set of spectral outputs. When in the overparameterised regime, the regression can harness this diverse response to produce a lower MSE in comparison with the spectrally limited WM and MS arrays.
Moving to parallel networks, we see that all combinations of parallel networks benefit from an overparameterised regime. This is expected when considering the effective dimensionality of the parallel network outputs. Each of the three networks has a distinct set of internal dynamics and output spectra, as such, combining them in parallel enhances the spectral diversity, enabling networks to benefit from an overparameterised regime. Comparing the parallel networks with different architectures, we see that the improvements gained depend on which networks are included. Combining the two low non-linearity reservoirs (MS+WM) shows a weaker improvement than combining creating a 2 parallel network with the non-linear PW sample as expected. For series networks, the level of improvements depend on the arrangement of nodes as seen in the underparameterised regime. WM->MS shows worse overparameterised performance than MSWM despite the theoretical internal dynamics being the same for each network. As we are assessing performance on a prediction task, both the separated node dynamics and combined dynamics of the system plays a role. MSWM is more capable of transferring previous information and increasing memory capacity, leading to better predictions and a lower MSE. As such, both the range of dynamics of each node, and there arrangement play a role. Interestingly, WMMS displays better improvements than when combining MS and WM in parallel. Here, the internal dynamics are the same and the difference arises from the input that the second layer receives. By feeding the output of one array to the input of the next, an additional set of diverse dynamics arises from the input sequence.
[…]